You invest 10000$ in a savings account today at j1 = 5%. you withdraw 500$ at the end of year 1, 550$ at the end of year 2, 600$ at the end of year 3, and so on (increasing the size of the annual withdrawal by 50$ each year with respect to the previous year). What is the account balance at the end of year 10?
Answer #1

rate positively..

Computation of balance at the end of 10th year
i ii=i*5% iii iv=i+ii-iii
year Beginning balance Interest earned Amount with drawn ending balance
0 10,000.00
1 10,000.00     500.00     500.00 10,000.00
2 10,000.00     500.00     550.00     9,950.00
3     9,950.00     497.50     600.00     9,847.50
4     9,847.50     492.38     650.00     9,689.88
5     9,689.88     484.49     700.00     9,474.37
6     9,474.37     473.72     750.00     9,198.09
7     9,198.09     459.90     800.00     8,857.99
8     8,857.99     442.90     850.00     8,450.89
9     8,450.89     422.54     900.00     7,973.44
10     7,973.44     398.67     950.00     7,422.11
Ans =     7,422.11